OpenStudy (anonymous):

Solve the differential equation dy/dx = (1+y)/(2+x).

OpenStudy (anonymous):

its a separable differential equation

OpenStudy (anonymous):

yes thats what we are studdying at the moment. but i got a mental block with this on - i dont know why>

OpenStudy (ash2326):

We have \[\frac{dy}{dx}= \frac{(1+y)}{(2+x)}\] It can be rearranged as \[\frac{dy}{1+y}=\frac{dx}{2+x}\] Now integrate both sides \[\int \frac{dy}{1+y}=\int \frac{dx}{2+x}\] Can you do this?

OpenStudy (anonymous):

yes ln (1 + y) = ln(2 + x) + C

OpenStudy (ash2326):

great, you are correct:)

OpenStudy (anonymous):

i'm still confused about how you rearranged the equation initially

OpenStudy (anonymous):

multiply by dx, then divide by 1/y+1

OpenStudy (ash2326):